OpenStudy (anonymous):

simplify the expression ∛(x^2/y^5 )

OpenStudy (anonymous):

\[\large{\frac{x^{\frac{2}{3}}}{y^{\frac{5}{3}}}}\]

OpenStudy (anonymous):

nope not the answer in bk of book

OpenStudy (anonymous):

(x^2/3)/(y^5/3)